OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 1, -6>; v = <6, 7, -3>; w = <-7, 5, 2>

OpenStudy (anonymous):

\[u\cdot v = u_1 v_1 + u_2 v_2 + \cdot\cdot\cdot + u_n v_n\]

OpenStudy (anonymous):

I only have -42, 35, and -6. After that, I do not know what to do really.

OpenStudy (anonymous):

just add

OpenStudy (anonymous):

so, it's -13?

OpenStudy (anonymous):

looks that way

OpenStudy (anonymous):

Okay ! Thank You :)

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

My apologies, are you positive? Is that how you really solve this, it just seems too simple.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Dot_product

OpenStudy (anonymous):

Pardon me, I guess it is that simple. :)

OpenStudy (anonymous):

lol